# Nonparametric bounds for vaccine effects in randomized trials

### v1

@@ 4.1 LP-based bounds
We derive bounds for each VE by solving an optimization problem that maximizes (upper bound) or minimizes (lower bound) the VE under constraints derived from the observed data distribution. To obtain sharp and valid bounds, as discussed later, we do not optimize the VE expression directly but instead its building blocks, namely $E_{IV}(Y^{a,m})=E_{VI}(Y^{a,m})=E(Y^{a,m})$. To begin with, we show how both the target parameters and the constraints can be expressed as linear functions of the $q_{i.ab}$, thereby framing the problem as an LP problem.
@@
where the first equality follows from randomization of $A$ and the last one by the definition of the potential response types. The constraints are derived from the observed conditional probability $p_{y.ab}$. As in the last line of Equation (7), $p_{y.ab}$ can be represented as the sum $p_{y.ab}=\sum_{i\in Q_{b}(a,m=-1)}q_{i.ab}$. Additional constraints are induced from $q_{i.ab}$ being probabilities that sum to one over the potential response types. Putting it together, for every combination of $A=a$ and $B=b$, the constraints are
$\begin{split}1.\,&p_{y.ab}=\sum_{i\in Q_{b}(a,m=-1)}q_{i.ab}\;\text{for }\;y\in\{0,1\}\\ 2.\,&\sum_{i=0}^{15}q_{i.ab}=1\\ 3.\,&q_{i.ab}\geq 0,\text{for}\,i=0,...,15\end{split}$ (8)
@@
From Equation (7), we have that
$VE(0)=1-\frac{\pi_{0.1}\sum_{i\in Q_{1,0}}q_{i.10}+\pi_{1.1}\sum_{i\in Q_{1,1}}q_{i.11}}{\pi_{0.0}\sum_{i\in Q_{0,0}}q_{i.00}+\pi_{1.0}\sum_{i\in Q_{0,1}}q_{i.01}},$
the numerator and denominator involve disjoint sets of the parameters $q_{i.ab}$. This separable structure implies $\min_{q}\frac{N(q)}{D(q)}=\frac{\min_{q}N(q)}{\max_{q}D(q)}$ and $\max_{q}\frac{N(q)}{D(q)}=\frac{\max_{q}N(q)}{\min_{q}D(q)}$. Therefore, optimizing the numerator and denominator separately yields valid bounds for $VE(0)$, which in this case are also sharp. The same separability applies to $VE(1)$ and $VE_{T}$, so their bounds can be obtained in the same way. In contrast, for $VE_{M}(a)$ the numerator and denominator share parameters $q_{i.ab}$, so this simplification is not available. As described in Appendix D, we solve that case using direct nonlinear optimization, which produces valid bounds but without a symbolic expression.
Finally, to solve an LP problem, Balke and Pearl, (1994) developed a program that takes symbolic LP expressions as input and returns symbolic solutions for the bounds, unlike standard LP solvers that provide only numerical values. Following Gabriel et al., 2023b , we instead used the rcdd R package (Geyer and Glen D Meeden,, 2023) for each of the four LP problems. The rcdd R package is more general and suitable for complex scenarios beyond those handled by the original program.
A full description of the LP method, together with a detailed derivation of the following result, is given in Appendix C.
@@
Consider the setting of Figure 2. The following bounds for $VE(m)$, $m\in\{0,1\}$ and $VE_{T}$ are valid and sharp.
@@ 4.2 Monotonicity-based bounds
In this section, we derive monotonicity-based bounds based on two monotonicity assumptions, different from those just described. The first assumes that both the belief $B$ and the outcome $Y$ are monotonically related to the unmeasured common cause $U$.
@@
[…]
  1. (i)
@@
An example of a variable $U$ that might satisfy Assumption 7(i) is optimism. Optimistic participants may engage in “wishful thinking” and believe they received the vaccine even when they did not (Bang,, 2016). They may also be more outgoing, thereby increasing their risk of infection. By contrast, pessimism provides an example of the same monotonicity but in the opposite direction: both $\Pr(Y=1\mid A=a,B=1,U=u)$ and $\Pr(B=1\mid A=a,U=u)$ decrease with $u$. An illustration of Assumption 7(ii) is neuroticism. Neurotic participants may be more anxious and thus more likely to believe they received the vaccine, consistent with findings that placebo effects are more common among them (Kern et al.,, 2020). Such participants often have more introverted personalities and may therefore avoid or minimize social contact, which reduces their risk of infection (Glei and Weinstein,, 2023).
The following proposition builds on Assumption 7 to provide nonparametric bounds for $VE(0)$ and $VE_{T}$.
Proposition 2.
Consider the setting of Figure 2. Under Assumptions 2 and 7(i), the following bounds for $VE(0)$ and $VE_{T}$ are valid and sharp.
$1-\frac{1}{p_{1.00}}$ $\leq VE(0)\leq 1-p_{1.10},$
$1-\frac{p_{1.11}}{p_{1.00}}$ $\leq VE_{T}\leq 1.$
Under Assumptions 2 and 7 (ii), the bounds for $VE(0)$ and $VE_{T}$ given in Appendix G are valid and sharp.
The proof is given in Appendix G. Unfortunately, Assumption 7 and the resulting Proposition 2 do not provide information on $VE(1)$ on an upper bound of $VE_{T}$. That is, without further assumptions, we cannot construct bounds more informative than the trivial bounds $\infty\leq VE(1)\leq 1$ and $VE_{T}\leq 1$ (the lower bound is given by Proposition 2). To remedy this problem, the second assumption we consider asserts that the sign of the effect of the message $M$ on the mean outcome $Y$ is known.
@@
This assumption is plausible under Assumptions 1 and 6. For instance, if a participant receives a message indicating vaccination ($M=1$), then by Assumption 1 this also implies $B=1$. Assumption 8(i) reflects the scenario in which believing that one is vaccinated is likely to increase the risk of infection by encouraging riskier behaviors, such as more frequent contacts or reduced adherence to safety guidelines (Trogen and Caplan,, 2021; Buckell et al.,, 2023). By contrast, Assumption 8(ii) reflects the opposite scenario, where believing one is vaccinated would reduce the risk of infection, a pattern that appears unlikely in practice.
@@
Consider the setting of Figure 2. Under Assumptions 2,7(i), and 8(i), the following bounds for $VE(m),\;m\in\{0,1\}$ and $VE_{T}$ are valid and sharp.
@@
Under Assumptions 7(ii) and 8(ii) the bounds for $VE(m)$, $m\in\{0,1\}$ and $VE_{T}$ given in Appendix H are valid and sharp.
The proof is given in Appendix H.
@@ 5 Bounds for vaccine effects under Figure 3(a)
Similarly to the scenario illustrated in Figure 2, Assumption 3 is also violated in Figure 3(a) due to the path $B\leftarrow U\rightarrow Y$. However, unlike Figure 2, Figure 3(a) introduces an additional factor: AEs ($S$) are observed and directly influence the belief ($B$). In vaccine trials, it is more realistic to assume that treatment ($A$) affects belief through AEs rather than directly. Consistent with this assumption, empirical evidence shows that participants in vaccine trials often rely on perceived AEs to infer their treatment allocation, leading to unblinding (Hróbjartsson et al.,, 2014).
[…]
@@ 5.1 LP-based bounds
[…]
$\begin{split}1.\,&p_{y.asb}=\sum_{i\in Q_{a,s,b}(a,m=-1)}q_{i.sab}\;\text{for }\;y\in\{0,1\}\\ 2.\,&\sum_{i=0}^{15}q_{i.asb}=1\\ 3.\,&q_{i.asb}\geq 0,\text{for}\,i=0,...,15\end{split}$ (11)
Now that we have expressed both the target parameter and the constraints in terms of $q_{i.asb}$, we can formulate the corresponding LP problem. Since Equations (10) and (11) must hold for both $s=0$ and $s=1$, the LP can be solved separately within each stratum of $s$. We then select, for the lower and upper bound separately, the value of $s$ that yields the tightest result. This procedure ensures that the bounds presented in next proposition are as sharp as possible.
@@
Consider the setting of Figure 3(a). The following bounds for the $VE(m)$, $m\in\{0,1\}$ and $VE_{T}$ are valid and sharp.
@@
The proof is given in Appendix I. The LP-based bounds in Proposition 4 are guaranteed to be at least as sharp as those in Proposition 1, since they are obtained by additionally conditioning on $S$ and then taking the minimum or maximum within each stratum.
@@ 5.2 Monotonicity-based bounds
Because under Figure 3(a) we have $S\perp_{\mathcal{T}_{VI}}Y\mid A,B,M,U$, we can replace Assumption 7 with the following requirement that conditions on $S$ and imposes monotonicity only within each stratum of $S$.
@@
[…]
  1. (i)
@@
As formally presented in the following Proposition, under Assumption 9, the expressions $p_{1.ab}$ in Propositions 2 and 3 can be replaced by $\min_{s\in\{0,1\}}\{p_{1.abs}\}$ or by $\max_{s\in\{0,1\}}=\{p_{1.abs}\}$ depending on whether they appear in the numerator or denominator and whether we are computing an upper or lower bound.
@@
Consider the setting of Figure 3(a). Under Assumptions 5, 8(i), and 9(i), the following bounds for $VE(m)$, $m\in\{0,1\}$ and $VE_{T}$ are valid and sharp.
@@
Under Assumptions 8(ii) and 9(ii), the bounds for $VE(m)$, $m\in\{0,1\}$ and $VE_{T}$ given in Appendix J are valid and sharp.
The proof is given in Appendix J. The monotonicity-based bounds in Proposition 5 are guaranteed to be at least as sharp as those in Proposition 3, since they are obtained by additionally conditioning on $S$. As in the LP case, this conditioning allows the bounds to be computed within each stratum of $S$, and then combined by taking the appropriate minimum or maximum, which can only tighten the results.
6 Bounds for vaccine effects under Figures 3(b), 3(c) and 3(d)
In this section, we discuss Figures 3(b), 3(c) and 3(d), where $S$ is measured. As we will show, however, $S$ cannot be used to sharpen the bounds in Propositions 1 and 3. We begin by describing how the role of AEs differs across these figures.
@@ 6 Bounds for vaccine effects under Figures 3(b), 3(c) and 3(d)
Figures 3(c) and 3(d) violate Assumption 4 due to the path $B\leftarrow U\rightarrow Y$. However, unlike Figure 3(b), they also include the path $S\leftarrow U\rightarrow Y$. Such scenario is possible, for example, if personality traits $U$ influence not only $B$ and $Y$ but also the perception and reporting of AEs ($S$). For instance, anxiety sensitivity and neuroticism have been linked to a heightened tendency to interpret mild symptoms as severe side effects, a phenomenon known as the nocebo effect (Meijers et al.,, 2022; Witthöft and Hiller,, 2022).
@@ 6.1 LP-based bounds
The bounds for the VE given in Proposition 1 are valid and sharp in the setting of Figures 3(b)3(d).
The proof is given in Appendix K.
@@ 6.2 Monotonicity-based bounds
As in the LP-based case, the sharp and valid bounds under Figures 3(b)3(d), together with Assumptions 5, 8 and 9, coincide with the bounds presented in Proposition 3.
@@
Under Assumptions 5, 8 and 9, the bounds for the VE given in Proposition 3 are valid and sharp in the setting of Figures 3(b)3(d).
The proof is given in Appendix L.
To emphasize the findings from this section, if either (i) the arrow $S\rightarrow Y$ exists, or (ii) the arrow $U\rightarrow S$ exists (in addition of being a common cause of $B$ and $Y$), then $S$ cannot be used to improve the width of the bounds, and the resulting bounds (in Propositions 67) remain the same as the results without observing $S$ (Propositions 13). While $B$ is not a treatment variable, these finding are reminiscing of a demand that $S$ has a similar role to an instrumental variable (with respect to $B$). Furthermore, finding (ii) aligns with the results of Jonzon et al., (2025) (their Section 4.1), who similarly demonstrated that when the covariate is confounded by an unobserved variable, “covariate averaging does not improve the width of the bounds,” and the covariate-adjusted and unadjusted bounds coincide.
As a remark, in some applications it may be more realistic to assume distinct unmeasured variables, $U_{1}$ and $U_{2}$, rather than a single common cause (Figure M.1). For example, $U_{1}$ could represent personality traits that affect both belief and outcome, while $U_{2}$ could represent pre-existing health conditions, such as chronic inflammation or immunodeficiency, that increase both the likelihood of strong side effects and susceptibility to infection (Polack et al.,, 2020). As shown in Appendix M, however, introducing $U_{2}$ does not affect the key independence relations and therefore does not alter the resulting bounds: $Y\perp_{\mathcal{T}_{VI}}M\mid A,B,S,U_{1}$ still holds, while $Y^{a,m}\not\perp_{\mathcal{T}_{VI}}S$, so Propositions 6 and 7 continue to apply.

### v2

@@ 4.1 LP-based bounds
We derive bounds for each VE by solving an optimization problem that maximizes (upper bound) or minimizes (lower bound) the VE under constraints derived from the observed data distribution. To obtain the bounds, as discussed later, we do not optimize the VE directly but instead its building blocks, namely $E_{IV}(Y^{a,m})=E_{VI}(Y^{a,m})=E(Y^{a,m})$. To begin with, we show how both the target parameters and the constraints can be expressed as linear functions of the $q_{i.ab}$, thereby framing the problem as an LP problem.
@@
where the first equality follows from randomization of $A$ and the last one by the definition of the potential response types. The constraints are derived from the observed conditional probability $p_{y.ab}$. As in the last line of Equation (7), $p_{y.ab}$ can be represented as the sum $p_{y.ab}=\sum_{i\in Q_{a,b}(a,m=-1)}q_{i.ab}$. Additional constraints are induced from $q_{i.ab}$ being probabilities that sum to one over the potential response types. Putting it together, for every combination of $A=a$ and $B=b$, the constraints are
$\begin{split}1.\,&p_{y.ab}=\sum_{i\in Q_{a,b}(a,m=-1)}q_{i.ab}\;\text{for }\;y\in\{0,1\}\\ 2.\,&\sum_{i=0}^{15}q_{i.ab}=1\\ 3.\,&q_{i.ab}\geq 0,\text{for}\,i=0,...,15\end{split}$ (8)
@@
Finally, each LP problem can be solved using the program developed by Balke and Pearl (1994), which takes symbolic LP expressions as input and returns symbolic solutions for the bounds, unlike standard LP solvers that provide only numerical values. Following Gabriel et al. (2023b), we instead use the rcdd R package (Geyer and Meeden, 2023) to solve each of the four LP problems. The rcdd R package is more general and suitable for complex scenarios beyond those handled by the original program. The following proposition presents the LP-based bounds under Figure 2..
@@
Consider the setting of Figure 2. The following bounds for $VE(m)$, $m\in\{0,1\}$ and $VE_{T}$ hold.
@@ 4.2 Monotonicity-based bounds
In this section, we derive monotonicity-based bounds based on two monotonicity assumptions, different from those just described. The first assumes that both the belief and the outcome are monotonically related to the unmeasured common cause.
@@
(U monotonicity) In the setting of Figure 2, suppose that $Y\perp_{VI}M|A,B,U$ and one of the following conditions hold:
  1. 1.
@@
An example of a variable $U$ that might satisfy Assumption 7.1 is optimism. Optimistic participants may engage in “wishful thinking” and believe they received the vaccine even when they did not (Bang, 2016). They may also be more outgoing, thereby increasing their risk of infection. By contrast, pessimism provides an example of the same monotonicity but in the opposite direction: both $\Pr(Y=1\mid A=a,B=1,U=u)$ and $\Pr(B=1\mid A=a,U=u)$ decrease with $u$. An illustration of Assumption 7.2 is neuroticism. Neurotic participants may be more anxious and thus more likely to believe they received the vaccine, consistent with findings that placebo effects are more common among them (Kern et al., 2020). Such participants often have more introverted personalities and may therefore avoid or minimize social contact, which reduces their risk of infection (Glei and Weinstein, 2023).
Unfortunately, under Assumption 7, we obtain two-sided bounds only for $VE(0)$. The resulting bounds do not provide information on $VE(1)$ or on an upper bound for $VE_{T}$ (see Appendix G). To remedy this problem, the second assumption we consider asserts that the sign of the effect of the message $M$ on the mean outcome $Y$ is known.
@@
This assumption is plausible under Assumptions 1 and 6. For instance, if a participant receives a message indicating vaccination ($M=1$), then by Assumption 1 this also implies $B=1$. Assumption 8.1 reflects the scenario in which believing that one is vaccinated is likely to increase the risk of infection by encouraging riskier behaviors, such as more frequent contacts or reduced adherence to safety guidelines (Trogen and Caplan, 2021; Buckell et al., 2023). By contrast, Assumption 8.2 reflects the opposite scenario, where believing one is vaccinated would reduce the risk of infection, a pattern that appears unlikely in practice.
@@
Consider the setting of Figure 2. Under Assumptions 2,7.1, and 8.1, the following bounds for $VE(m),\;m\in\{0,1\}$ and $VE_{T}$ hold.
@@
Under Assumptions 2, 7.2 and 8.2 the bounds for $VE(m)$, $m\in\{0,1\}$ and $VE_{T}$ given in Appendix H hold.
@@ 5 Bounds for vaccine effects under Figure 3(a)
Similarly to the scenario illustrated in Figure 2, Assumption 3 is also violated in Figure 3(a) due to the path $B\leftarrow U\rightarrow Y$. However, unlike Figure 2, Figure 3(a) introduces an additional factor: AEs ($S$) are observed and directly influence the belief ($B$). In vaccine trials, it is more realistic to assume that treatment ($A$) affects belief through AEs, rather than through a direct effect on belief. Consistent with this assumption, empirical evidence shows that participants in vaccine trials often rely on perceived AEs to infer their treatment allocation, leading to unblinding (Hróbjartsson et al., 2014).
Under Figure 3(a), the arrows $U\to S$ and $S\to Y$ are absent. This implies two additional independence assumptions: $S\perp_{\mathcal{T}_{VI}}U$ and $Y\perp_{\mathcal{T}_{VI}}S\mid A,B,M,U$. As we show in this section, these two independence assumptions are exactly what allow us to obtain narrower bounds under both the LP-based and the monotonicity-based approaches.
@@ 5.1 LP-based bounds
where the first equality follows from the randomization of $A$ and because $A$ blocks the only backdoor path between $S$ and $Y$ under Figure 3(a), so $Y^{a,m}\perp_{\mathcal{T}_{VI}}A,S^{a,m=-1}$ (for a proof see Appendix B.2). Moving to the constraints, similarly to Equation (8), we obtain the following system of linear constraints for every combination of $A=a$, $S=s$ and $B=b$:
$\begin{split}1.\,&p_{y.asb}=\sum_{i\in Q_{a,s,b}(a,m=-1)}q_{i.asb}\;\text{for }\;y\in\{0,1\}\\ 2.\,&\sum_{i=0}^{15}q_{i.asb}=1\\ 3.\,&q_{i.asb}\geq 0,\text{for}\,i=0,...,15\end{split}$ (11)
Now that we have expressed both the target parameter and the constraints in terms of $q_{i.asb}$, we can formulate the corresponding LP problem. Since Equations (10) and (11) must hold for both $s=0$ and $s=1$, the LP can be solved separately within each stratum of $s$. We then select, for the lower and upper bound separately, the value of $s$ that yields the tightest result. This procedure ensures that the bounds presented in the next proposition are narrower than the bounds in Proposition 1.
@@
Consider the setting of Figure 3(a). The following bounds for the $VE(m)$, $m\in\{0,1\}$ and $VE_{T}$ hold.
@@ 5.2 Monotonicity-based bounds
Because under Figure 3(a) we have $S\perp_{\mathcal{T}_{VI}}Y\mid A,B,M,U$, we can replace Assumption 7 with the following requirement that conditions on $S$ and imposes monotonicity within each stratum of $S$.
@@
($U,S$ monotonicity) Consider the setting of Figure 3(a). Suppose that $Y\perp_{\mathcal{T}_{VI}}M|A,B,S,U$ and one of the following conditions holds:
  1. 1.
@@
As formally presented in the following Proposition, under Assumption 9, the expressions $p_{1.ab}$ in Proposition 2 can be replaced by $\min_{s\in\{0,1\}}\{p_{1.abs}\}$ or by $\max_{s\in\{0,1\}}=\{p_{1.abs}\}$ depending on whether they appear in the numerator or denominator and whether we are computing an upper or lower bound.
@@
Consider the setting of Figure 3(a). Under Assumptions 5, 8.1, and 9.1, the following bounds for $VE(m)$, $m\in\{0,1\}$ and $VE_{T}$ hold.
@@
Under Assumptions 5, 8.2 and 9.2, the bounds for $VE(m)$, $m\in\{0,1\}$ and $VE_{T}$ given in Appendix J hold.
The monotonicity-based bounds in Proposition 4 are guaranteed to be at least as narrow as those in Proposition 2, since they are obtained by additionally conditioning on $S$. As in the LP case, this conditioning allows the bounds to be computed within each stratum of $S$, and then combined by taking the appropriate minimum or maximum, which can only tighten the results.
6 Bounds for vaccine effects under Figures 3(b)3(d)
In this section, we discuss Figures 3(b)3(d), where $S$ is measured. As we will show, however, $S$ cannot be used to obtain narrower bounds in Propositions 1 and 2. We begin by describing how the role of AEs differs across these figures.
@@ 6 Bounds for vaccine effects under Figures 3(b)3(d)
Figures 3(c)3(d) violate Assumption 4 due to the path $B\leftarrow U\rightarrow Y$. However, unlike Figure 3(b), they also include the path $S\leftarrow U\rightarrow Y$. Such scenario is possible, for example, if $U$ influence not only $B$ and $Y$ but also the perception and reporting of AEs ($S$). For instance, anxiety sensitivity and neuroticism have been linked to a heightened tendency to interpret mild symptoms as severe AEs, a phenomenon known as the nocebo effect (Meijers et al., 2022; Witthöft and Hiller, 2022).
@@ 6.1 LP-based bounds
The bounds for the VE given in Proposition 1 hold in the setting of Figures 3(b)3(d).
@@ 6.2 Monotonicity-based bounds
In the case of Figures 3(b)3(d), to obtain bounds, in addition to the conditions in Assumption 9, we also assume the directions of the effects of $S$ on $B$ and of $S$ on the conditional expectation $E(Y^{a,m}\mid A=a,S=s)$.
Assumption 10.
($U|S$ monotonicity) Consider the setting of Figures 3(b)3(d). Suppose that $Y\perp_{\mathcal{T}_{VI}}M|A,B,S,U$, and the following conditions hold:
  1. 1.
  2. 2.
    $\Pr(B=1|A=a,S=s)$ is nondecreasing in $s$, for all $a$.
  3. 3.
    $E(Y^{a,m}|A=a,S=s)$ is nondecreasing in $s$, for all $a$.
Next, we clarify the additional conditions in Assumption 10. Condition 10.2 captures the idea that individuals who experience AEs are more likely to believe that they received the vaccine, which is typically plausible in vaccine trials. To explain the meaning of Condition 10.3, we consider Figure 3(c), in which the path $S\to Y$ is absent. Under this figure, Condition 10.3 is implied by two simpler monotonicity assumptions: (i) $\Pr(Y=1\mid A=a,B=1,U=u)$ is nondecreasing in $u$; and (ii) $\Pr(S=1\mid U=u)$ is nondecreasing in $u$. Intuitively, Condition 10.3 means that both the outcome and the occurrence of AEs increase with $U$, so higher values of $S$ correspond to higher values of $E(Y^{a,m})$. Nevertheless, since the arrow $S\to Y$ may exist in general, the simplified monotonicity conditions serve only as intuition, and the analysis still requires Condition 10.3 itself.
As in the LP-based case, the bounds under Figures 3(b)3(d), together with Assumptions 5, 8 and 10, coincide with the bounds presented in Proposition 2.
@@
Under Assumptions 5, 8 and 10, the bounds for the VE given in Proposition 2 hold in the setting of Figures 3(b)3(d).
The assumption corresponding to Assumption 10 that yields reversed bounds is provided in Appendix L.4.
To emphasize the findings from this section, if either (i) the arrow $S\rightarrow Y$ exists, or (ii) the arrow $U\rightarrow S$ exists (in addition of being a common cause of $B$ and $Y$), then $S$ cannot be used to improve the width of the bounds, and the resulting bounds (in Propositions 56) remain the same as the results without observing $S$ (Propositions 12). While $B$ is not a treatment variable, these finding are reminiscing of a demand that $S$ has a similar role to an instrumental variable (with respect to $B$). Furthermore, finding (ii) aligns with the results of Jonzon et al. (2025) (their Section 4.1), who similarly demonstrated that when the covariate is confounded by an unobserved variable, “covariate averaging does not improve the width of the bounds,” and the covariate-adjusted and unadjusted bounds coincide.
As a remark, in some applications it may be more realistic to assume distinct unmeasured variables, $U_{1}$ and $U_{2}$, rather than a single common cause (Figure M.1). For example, $U_{1}$ could represent personality traits that affect both belief and outcome, while $U_{2}$ could represent pre-existing health conditions, such as chronic inflammation or immunodeficiency, that increase both the likelihood of strong AEs and susceptibility to infection (Polack et al., 2020). As shown in Appendix M, however, introducing $U_{2}$ does not affect the key independence relations and therefore does not alter the resulting bounds: $Y\perp_{\mathcal{T}_{VI}}M\mid A,B,S,U_{1}$ still holds, while $Y^{a,m}\not\perp_{\mathcal{T}_{VI}}S$, so Propositions 5 and 6 continue to apply.